You have just read that chickens have a relatively low level of body fat. The population mean percentage of fat in a chicken is reported to be 32% with the standard deviation of 8.7%. How likely is it that you could randomly select a sample of 12 chickens from the population which would have a mean of 26% body fat or less?
Answer #1

Given ,

= 32 , = 8.7

using central limit theorem,

P( < x) = P(Z < ( x - ) / ( / sqrt(n) ) )

So,

P( <= 26) = P(Z <= ( 26 - 32) / ( 8.7 / sqrt(12) ) )

= P(Z <= -2.39)

= .0.0084